1/3

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

Der Werwolf La lupfanto- mo The Werewolf The Banshee (An Approach)

tradukita de Bertilo Wennergren tradukita de Alexander Gross tradukita de Max Knight

Ein Werwolf eines

Nachts entwich von Weib und Kind

und sich begab an eines Dorfschul-

lehrers Grab und bat ihn: �Bitte,

beuge mich!“ Jen lupfantom’ de famili’ meznokte iam kuris for al tombo de la profesor’. kaj petis pri konjugaci’.

A Werewolf, troubled by

his name, Left wife and brood one night and came To a hidden graveyard to

enlist The aid of a long-dead phi-

lologist.

One night, a banshee slunk away from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect

me, pray.”

Der Dorfschulmeister

stieg hinauf auf seines Blech-

schilds Messingknauf

und sprach zum Wolf,

der seine Pfoten geduldig kreuzte vor

dem Toten:

Do diktis tiu dok-

ta ul’ starante sur la kructabul’ dum lup’ senmova kiel splinto atentis vortojn de l’ mortinto.

“Oh sage, wake up, please don’t berate me,” He howled sadly, “Just con-

jugate me.” The seer arose a bit unstea-

dy Yawned twice, wheezed once, and then was ready.

The village teacher climbed up straight upon his grave stone with its plate and to the apparition

said who meekly knelt before the dead:

�Der Werwolf“ – sprach der gute Mann,

�des Werwolfs, Geni-

tiv sodann, dem Werwolf, Dativ,

wie man’s nennt, den Werwolf, - damit

hat’s ein End!“ De “lupfantomo”, simpla rekta, tra “lupfintomo”, plej perfekta, plu “lupfontomo”, iom rara, ĝis “lupfuntomo”, ho, koŝmara.

“Well, ‘Werewolf’ is your

plural past, While ‘Waswolf’ is singu-

larly cast: There’s ‘Amwolf’ too, the

present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

Dem Werwolf schmei-

chelten die Fälle er rollte seine Au-

genbälle. �Indessen“, bat er,

�füge doch zur Einzahl auch die

Mehrzahl noch!“ Ĉi participoj plaĉis multe al la fantomo, kaj rezulte ĝi petis: “Legu plu el PIV-o pri participoj en pasivo.”

“I know that–I’m no men-

tal cripple– The future form and parti-

ciple Are what I crave,” the be-

ast replied. The scholar paused–again

he tried:

The banshee marveled at the cases and writhed with pleasure, making faces, but said: “You did not add, so far, the plural to the singu-

lar!”

…  …  …  …

2/3

Der Dorfschullehrer

aber musste gestehn, dass er von

ihr nichts wusste. Zwar Wölfe gab’s in

großer Schar, doch �Wer“ gab’s nur

im Sinular.

La viro ĉagreniĝis tamen: “Jam fermis mi per fin’ kaj amen, ĉar netransiraj bestoj feblas kaj la pasiv’ neniam eblas.”

“A ‘Will-be-wolf?’ It’s just

too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond all suture– I’m sorry, but you have no

future.”

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be pluri-

zed.”

Der Wolf erhob sich

tränenblind – er hatte ja doch Weib

und Kind! Doch da er keine Ge-

lehrter eben, so schied er dankend

und ergeben.

Sed lupfantomo pri transiro ja spertas pli ol eĉ vampiro. La lup’ revenis larmokula al sia famili’ ulula.

The Werewolf knew better–his sons still slept

At home, and homewards

now he crept, Happy, humble, without

apology For such folly of philology.

The banshee, rising

clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

…  …  …  …

3/3

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en Esperanton de Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu  la  re-

tejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .